OpenStudy (anonymous):

I WILL MEDAL AND FAN!!! ANSWERS ASAP PLEASE! What is the outlier of 4, 4, -6, -2, 14, 1, 1? A. 14; it RAISES the mean by about 1.9. B. 14; it LOWERS the mean by about 1.9.

OpenStudy (anonymous):

PLEASE HELP!!

OpenStudy (rizags):

definitely A

OpenStudy (anonymous):

I know it is 14, I eliminated 12 and 10 for the answer choices.

OpenStudy (anonymous):

Oh ok!

OpenStudy (anonymous):

Any evidence proof?

OpenStudy (anonymous):

if you give proof i will medal AND fan

OpenStudy (rizags):

adding a 14 to that set (which is bigger than all of the others) MUST increase the mean, i'll prove it here:

OpenStudy (rizags):

Without 14 : (4+4-6-2+1+1)/6 = 1/3, or about 0.33333 With 14: (4+4-6-2+1+1+14)/7 = 16/7, or about 2.29 their difference is about 1.9, with the addition of 19 INCREASING the original mean

OpenStudy (rizags):

i mean the addition of 14

OpenStudy (anonymous):

THANKS!! I fan and medaled :D thanks sooooo much dude!

OpenStudy (rizags):

ye np